MONT 106Q – Mathematical Thinking Third Writing Assignment

The third of the larger paper assignments this semester will be devoted to the pairing Shakespeare’s Othello and Mehmet Ali Sanlikol’s Othello in the Seraglio. For this assignment I want you to think about the relationship between these two versions of the basic story, the choices Sanlikol made in constructing his “coffee-house opera,” and whether this alternate version adds an interesting sidelight to the story. This will mostly involve mostly discussing your personal reactions, so it will not be as formal as the previous paper and there is definitely more room for your opinion and personal voice to come through.

Your essay should be about 2 pages in length (double-spaced, single-sided). Please submit via Google Docs or as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Monday, October 31.

Paper topic

First, there is a consistent “maneuver” that Sanlikol and his librettist (i.e. the writer of the text that is sung in the opera) Robert Labaree apply to the Othello story to construct the story of Sümbül the black eunuch in Othello in the Seraglio. What is this? Then, in your opinion, does this “maneuver” illuminate the many borders involved in this story, for instance the borders between Venice and Turkey, between white and black, between Muslim and non-Muslim? If so, how? If not, why not?

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. I’m going to be pretty much out of contact at a regional meeting of the American Mathematical Society out of town over the weekend of October 29 and 30, so I’m not going to be able to offer the usual comments on drafts this time.

2. Start by clearly stating your understanding of the “maneuver” alluded to in the topic description above in your first paragraph.

3. In the following paragraphs, by drawing on your understanding of the Shakespeare Othello, our class discussions, and your experience of the “coffee-house opera,” present your opinion about whether the play and the opera work together to create a deeper understanding of the boundaries involved in the story.

4. If you think they do, then explain how, by citing specific situations from the play and the opera, how they interact, and how one of them provides another point of view or a different slant on the other.

5. On the other hand, you’re certainly within your rights to say, “no, the play and the opera don’t connect in a meaningful way, or provide any illumination of each other, for me.” But if so, then it’s your job to explain why not. (It’s not sufficient just to state your opinion without providing the reasons that lead you to say that.)

6. You may present evidence consisting of quotations of passages from the Shakespeare play if you want, together with interpretations of those passages that come from your own thinking about them. Your evidence should be as specific, detailed, correct, and as varied as possible (focus on small details and mention the little things that support your point of view). If you do use quotations, identify the act, the scene number, and the line number(s) in the Folger Shakespeare Library edition of Othello.

7. Your essay should be well-organized as a whole and “flow” from one point to the next. Do not repeat pieces of evidence unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

8. As always, your sentences should be clear, readable, and not overly wordy. Aim for a reasonably formal but lively style. For this assignment, use of the first person is more or less expected while you are presenting your opinion about the effectiveness of the pairing of Othello with Othello in the Seraglio. Generally, colloquial ways of saying things that are acceptable in casual conversation are not as appropriate in this kind of writing, but if you have a really good reason for using a colloquial expression and it fits well, that’s OK. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

9. And as on previous assignments, errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.